OpenStudy (anonymous):

equations is of a parabola with a vertex at (0, 6)?

OpenStudy (imstuck):

Do you have the location of the focal point or the line of the directrix?

OpenStudy (anonymous):

0, 6

OpenStudy (imstuck):

That's the vertex. That alone does not tell me if the parabola is of the x^2 or y^2 in nature. You have nothing else but the vertex?

OpenStudy (imstuck):

Not the focus? Not the directrix?

OpenStudy (imstuck):

You cannot write this equation without either the focus or the directrix.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y = (x - 6)^2 y = (x + 6)^2 y = x^2 - 6 y = x^2 + 6

OpenStudy (anonymous):

its one of those

OpenStudy (imstuck):

It could look like either one of these, in general: if it's an x^2, which it appears to be.